a man when asked how many hens and buffaloes he has, told that his animals have 120 eyes and 180 legs. How many hens and buffaloes has he?
Answer by rwm(914)

hens and buffaloes both have two eyes each
hens have two legs
and buffaloes have 4 leg
4b+2h=180 legs
2h+2b=120 eyes
or
h+b=60 animals
30 of each
30*4=120
30*2=60
120+60=180 legs